# Imprints of an early matter-dominated era arising from dark matter dilution mechanism on cosmic string dynamics and gravitational wave signatures

### v2

@@ III.2 Dark matter dilution mechanism
In this subsection, we discuss the DM dilution mechanism as the origin of the EMD era. As severe constraints from DM detection experiments have reduced interest in weakly interacting massive particles with masses around the electroweak scale as DM candidates, a lighter DM candidate $X$ has gained prominence due to its ability to evade direct detection bounds. Nevertheless, thermal production of $X$ particles with low annihilation cross sections typically results in an overproduction problem [86, 87]. Within the DM dilution mechanism [60, 61, 62, 63, 64, 66, 67], the overproduction of DM particles is mitigated by entropy injection from the decays of a dilutor particle $Y$, which dominates the universe for a period, thereby inducing an EMD era.
@@
where $T$ denotes the temperature and $\zeta(x)$ represents the Riemann zeta function. For sufficiently weak interactions, $X$ particles freezes out from the plasma at a temperature $T_{\mathrm{FO}}\gg m_{X}$, leaving a number density in the comoving volume of
@@
where $s_{0}=2891.2~{}\mathrm{cm}^{-3}$ and $\rho_{\mathrm{c}}=1.05\times 10^{-5}h^{2}~{}\mathrm{GeV}/\mathrm{cm}^{3}$ [76] are the entropy density and the critical energy density at present, respectively. Thus, for $m_{X}\lesssim$\mathrm{k}\mathrm{e}\mathrm{V}$$, $\Omega^{0}_{X}$ exceeds the observed cold DM relic density $\Omega_{\mathrm{CDM}}=0.265$ [76] by at least one order of magnitude, causing an overproduction problem.
Introducing a long-lived dilutor $Y$ with a mass $m_{Y}$ much larger than $m_{X}$ can effectively address the overproduction problem of $X$ particles. First, during the RD era, both $Y$ and $X$ particles decouple relativistically at a similar temperature, resulting in comparable yields, $Y_{Y}\simeq Y_{X}$. Second, because of $m_{Y}\gg m_{X}$, $Y$ particles become nonrelativistic at a relatively high temperature, while $X$ particles remain relativistic. Consequently, $Y$ particles quickly dominate the energy density of the universe, initiating an EMD era. Finally, when the lifetime of $Y$ particles comes to an end, they decay into SM particles and $X$ particles, injecting entropy and consequently diluting the energy density $\rho_{X}$ of $X$ particles.
@@
In the minimal LRSM with the $\mathrm{SU}(3)_{\mathrm{C}}\times\mathrm{SU}(2)_{\mathrm{L}}\times\mathrm{SU}(2)_{\mathrm{R}}\times\mathrm{U}(1)_{B-L}$ gauge symmetry [88, 89] considered in Ref. [68], there are two scenarios for the DM dilution mechanism, where the DM candidate $X$ is the lightest right-handed neutrino $N_{1}$.
@@
    where the family indices are suppressed. $\Phi$ is an $\mathrm{SU}(2)_{\mathrm{L}}\times\mathrm{SU}(2)_{\mathrm{R}}$ bidoublet scalar, with $\tilde{\Phi}=\mathrm{i}\sigma_{2}\Phi^{\ast}\mathrm{i}\sigma_{2}$. $\Delta_{\mathrm{L}}$ and $\Delta_{\mathrm{R}}$ are $\mathrm{SU}(2)_{\mathrm{L}}$ and $\mathrm{SU}(2)_{\mathrm{R}}$ triplet scalars, respectively. $Q_{\mathrm{L},\mathrm{R}}$ and $L_{\mathrm{L},\mathrm{R}}$ denote left-handed and right-handed quarks and leptons. $Y_{q}$, $\tilde{Y}_{q}$, $Y_{l}$, $\tilde{Y}_{l}$, $Y_{\Delta_{\mathrm{L}}}$, and $Y_{\Delta_{\mathrm{R}}}$ represent the Yukawa coupling matrices. The dilutor $\Delta$ is the Higgs boson arising from the electrically neutral component of $\Delta_{\mathrm{R}}$. Its decay channels involve $\Delta\rightarrow N_{1}N_{1}$, $\Delta\rightarrow W_{\mathrm{R}}^{\ast}W_{\mathrm{R}}^{\ast}\to\text{4 fermions}$, $\Delta\rightarrow\gamma\gamma$, and so on.
@@
We solve the Boltzmann equations for the benchmark points (BPs) specified in Tables 1 and 2 for Scenarios 1 and 2, respectively. The values of $y$, $m_{N_{1}}$, $m_{N_{2}}$, $m_{\Delta}$, $m_{W_{\mathrm{R}}}$, the ratio $\tan\beta$ of two vacuum expectation values from the bidoublet, and the mixing angle $\theta_{\Delta h}$ between $\Delta$ and the $125~{}$\mathrm{G}\mathrm{e}\mathrm{V}$$ Higgs boson $h$ are adopted based on the results presented in Ref. [68]. The dilutor decay width $\Gamma_{Y}$ and the branching ratio $B_{X}$ are calculated using the formulas provided therein. All BPs are selected to ensure the correct DM relic abundance through the DM dilution mechanism. BP1a and BP2a will be the primary focus of the following analysis, while the remaining BPs will provide additional results for comparison.
We adopt the initial temperature as $T_{\mathrm{ini}}=m_{Y}/10$, at which $Y$ particles have decoupled with $Y_{Y}\simeq Y_{X}$ and become nonrelativistic, while $X$ particles are relativistic. The initial conditions for the energy densities of $Y$, $X$, and SM particles are
@@
By substituting the obtained $\rho_{Y}$, $\rho_{X}$, and $\rho_{\mathrm{SM}}$ into Eq. (45) and solving Eq. (9) with $a(t_{0})=1$, we determine the evolution of the scale factor $a(t)$ in both BPs, as illustrated in Fig. 4. Compared with the standard $\Lambda\mathrm{CDM}$ cosmological model, the presence of the EMD era reduces the scale factor prior to $t_{2}$. This occurs because the scale factor evolves as $a\propto t^{2/3}$ during an MD era, increasing more rapidly with time than $a\propto t^{1/2}$ during an RD era. Therefore, and a smaller $a$ is required at the onset of the EMD era to ensure $a(t_{0})=1$ at the present time $t_{0}$.
@@ III.3 Impact on the loop number density of cosmic strings
Since the insertion of the EMD era affects the evolution of the CS network in the VOS model, it is necessary to examine how the normalized correlation length $\xi$ and the RMS velocity $v$ change with time. We should consider these variations when calculating the number density of CS loops during the EMD era. To achieve this, we solve Eqs. (16) and (17) with the Hubble rate $H(t)$ modified by the presence of the EMD era, and obtain $\xi$ and $v$ being functions of cosmic time $t$ for BP1a and BP2a, as demonstrated by blue solid lines in Fig. 5. For comparison, the results of the standard $\Lambda\mathrm{CDM}$ model without the EMD era are also presented as red dotted lines.
@@
During the RD era in the $\Lambda\mathrm{CDM}$ model, $\xi$ and $v$ basically maintain the scaling values $\xi_{\mathrm{r}}=0.271$ and $v_{\mathrm{r}}=0.662$, as stated in Eq. (18). Some deviations around $t\sim 10^{20}~{}$\mathrm{G}\mathrm{e}\mathrm{V}^{-1}$$ and $t\sim 10^{26}~{}$\mathrm{G}\mathrm{e}\mathrm{V}^{-1}$$ are attributed to changes in relativistic degrees of freedom shown in Eq. (12). For $t\gtrsim t_{\mathrm{eq}}$, the universe transits to the MD era, resulting in an increase in $\xi$ and a decrease in $v$. It is important to note that the time axis in Fig. 5 is on a logarithmic scale, indicating that the changes in $\xi$ and $v$ occur relatively slowly. Nonetheless, this represents a mild nonscaling effect.
When the EMD era occurs, $\xi$ increases, while $v$ decreases, and both quantities gradually return to their scaling values after the EMD era ends. Therefore, the EMD era also introduces a nonscaling effect, which will be incorporated into the calculation of the CS loop number density. Furthermore, we estimate the magnitudes of the time derivatives $\dot{\xi}$ and $\dot{v}$ by calculating the evolution of their ratios to $\xi/t$ and $v/t$, respectively, as shown in Fig. 6. We find that, during the EMD era for both BPs, $|t\dot{\xi}/\xi|$ and $|t\dot{v}/v|$ are smaller than $0.15$ and $0.03$, respectively. Thus, $\dot{\xi}$ and $\dot{v}$ are negligible compared to $\xi/t$ and $v/t$, and this will simplify the calculations below.
@@
where $\Gamma G\mu\ll\alpha_{\mathrm{m}}\xi_{\star}=0.18$ has been used in the third step.
In addition, similar to Eq. (35), we must account for the number densities of CS loops produced in a preceding era and surviving into subsequent eras, with respect to the timeline in Fig. 2. For clarity, we define the notations for these loop number densities as follows.
@@
which is just Eq. (35). Similarly, by introducing $l_{1,2}=l+\Gamma G\mu(t-t_{1,2})$, we obtain
@@ III.4 Modification of the GW spectrum
In order to assess the sensitivity of GW detection experiments, we also show the constraints from the North American Nanohertz Observatory for Gravitational Waves (NANOGrav) [90], the European Pulsar Timing Array (EPTA) [91], and the Parkes Pulsar Timing Array (PPTA) [92], as well as the sensitivity curves of the International Pulsar Timing Array (IPTA) [93], the Square Kilometer Array (SKA) [94], the Laser Interferometer Space Antenna (LISA) [52], TianQin [95], Taiji [96], LIGO [97], the Cosmic Explorer (CE) [98], and the Einstein Telescope (ET) [99].
@@
Figure 7: GW spectra from cosmic strings with $G\mu=10^{-11}$ modified by the EMD era for BP1a in Scenario 1 (a) and BP2a in Scenario 2 (b). The red solid lines represent the total GW spectra, which include contributions from CS loops across all eras, while the blue, green, olive, and cyan dotted lines correspond to the contributions from CS loops formed during the original RD, EMD, second RD, and final MD eras, respectively. For comparison, various constraints and sensitivity curves from several GW detection experiments are also shown.
Compared to the GW spectrum for $G\mu=10^{-11}$ in the $\Lambda\mathrm{CDM}$ model illustrated in Fig. 1, the spectra in both BPs with the EMD era display a suppression at high frequencies $f\gtrsim 10~{}$\mathrm{H}\mathrm{z}$$, which corresponds to the contributions from CS loops formed during the original RD and EMD eras. The primary reason for this suppression is as follows. Since the scaling behavior of the CS network is only slightly violated, the correlation length $L$ remains approximately proportional to the cosmic time $t$, suggesting that the lengths of the generated loops are positively correlated with the scale factor $a(t)$. As shown in Fig. 4, the EMD era reduces the scale factor before $t_{2}$. This means that CS loops with a given initial length $l$, which is inversely proportional to the frequencies of the emitted GWs through $f_{\mathrm{e}}=2n/l$, are formed at a later time, when the energy densities of both CS loops and emitted GWs are reduced. As a result, the GW spectrum at sufficiently high frequencies is suppressed, and the end time of the EMD era, $t_{2}$, is related to the GW frequency at which this suppression effect becomes significant. For the parameters listed in Table 1, the dilutor $N_{2}$ in BP1a has a longer lifetime than the dilutor $\Delta$ in BP2a, resulting in a longer EMD era and, consequently, a stronger suppression.
@@
In addition, we consider the results for BP1b and BP1c in Scenario 1 and for BP2b and BP2c in Scenario 2, whose parameters as tabulated in Tables 1 and 2, respectively. The obtained GW spectra are illustrated in Fig. 9. For BP1a, BP1b, and BP1c in Scenario 1, the dilutor $N_{2}$ has the same mass, but the masses of the DM candidate $N_{1}$ differ, causing variations in the decay width of $N_{2}$. A smaller decay width $\Gamma_{N_{2}}$ corresponds to a longer duration of the EMD era, leading to stronger suppression effects at high frequencies, as illustrated in Fig. 9(a). On the other hand, the masses of the dilutor $\Delta$ are different for BP2a, BP2b, and BP2c in Scenario 2. A heavier $\Delta$ implies that the EMD era occurs earlier, and hence a higher frequency at which the suppression of the GW spectrum commences. This is clearly demonstrated in Fig. 9(b).
@@
where $[f_{\mathrm{min}},f_{\mathrm{max}}]$ is the accessible frequency band. If the SNR reaches a threshold $\varrho_{\mathrm{thr}}=10$, it is probable that the GW signal will be detected.
@@
For $G\mu\lesssim 10^{-12}$, we find that the inclusion of the EMD era has insignificant impact on the SNRs, because the suppression effect occurs at frequencies outside the sensitive bands of LIGO, CE, and ET (cf. Fig. 8). In contrast, for $G\mu\gtrsim 10^{-12}$, the SNRs are obviously reduced. Compared to BP2a, the dilutor in BP1a has a longer lifetime, leading to a stronger suppression effect and, consequently, smaller SNRs.

### v3

@@ III.2 Dark matter dilution mechanism
In this subsection, we discuss the DM dilution mechanism as the origin of the EMD era. As severe constraints from DM detection experiments have reduced interest in weakly interacting massive particles with masses around the electroweak scale as DM candidates, a lighter DM candidate $X$ has gained prominence due to its ability to evade direct detection bounds. Nevertheless, the thermal production of $X$ particles with low annihilation cross sections typically results in an overproduction problem [92, 93]. Within the DM dilution mechanism [60, 61, 62, 63, 64, 66, 67], the overproduction of DM particles is mitigated by entropy injection from the decays of a dilutor particle $Y$, which dominates the universe for a period, thereby inducing an EMD era.
@@
where $T$ denotes the temperature and $\zeta(x)$ represents the Riemann zeta function. For sufficiently weak interactions, $X$ particles freeze out from the plasma at a temperature $T_{\mathrm{FO}}\gg m_{X}$, leaving a number density in the comoving volume of
@@
where $s_{0}=2891.2~\mathrm{cm}^{-3}$ and $\rho_{\mathrm{c}}=1.05\times 10^{-5}h^{2}~\mathrm{GeV}/\mathrm{cm}^{3}$ [76] are the present time entropy density and the critical energy density, respectively. Thus, for $m_{X}\lesssim$\mathrm{k}\mathrm{e}\mathrm{V}$$, $\Omega^{0}_{X}$ exceeds the observed cold DM relic density $\Omega_{\mathrm{CDM}}=0.265$ [76] by at least one order of magnitude, causing an overproduction problem.
Introducing a long-lived dilutor $Y$ with a mass $m_{Y}$ much larger than $m_{X}$ can effectively address the problem of $X$ particle overproduction. First, during the RD era, both $Y$ and $X$ particles decouple relativistically at a similar temperature, resulting in comparable yields, $Y_{Y}\simeq Y_{X}$. Second, because $m_{Y}\gg m_{X}$, $Y$ particles become nonrelativistic at a relatively high temperature, while $X$ particles remain relativistic. Consequently, $Y$ particles quickly dominate the energy density of the universe, initiating an EMD era. Finally, when the lifetime of $Y$ particles comes to an ends, they decay into SM particles and $X$ particles, injecting entropy and consequently diluting the energy density $\rho_{X}$ of $X$ particles.
@@
In the minimal LRSM with the $\mathrm{SU}(3)_{\mathrm{C}}\times\mathrm{SU}(2)_{\mathrm{L}}\times\mathrm{SU}(2)_{\mathrm{R}}\times\mathrm{U}(1)_{B-L}$ gauge symmetry [94, 95] considered in Ref. [68], two scenarios exist for the DM dilution mechanism, where the DM candidate $X$ is the lightest right-handed neutrino $N_{1}$.
@@
    where the family indices are suppressed. $\Phi$ is an $\mathrm{SU}(2)_{\mathrm{L}}\times\mathrm{SU}(2)_{\mathrm{R}}$ bidoublet scalar, with $\tilde{\Phi}=\mathrm{i}\sigma_{2}\Phi^{\ast}\mathrm{i}\sigma_{2}$. $\Delta_{\mathrm{L}}$ and $\Delta_{\mathrm{R}}$ are $\mathrm{SU}(2)_{\mathrm{L}}$ and $\mathrm{SU}(2)_{\mathrm{R}}$ triplet scalars, respectively. $Q_{\mathrm{L},\mathrm{R}}$ and $L_{\mathrm{L},\mathrm{R}}$ denote left-handed and right-handed quarks and leptons. $Y_{q}$, $\tilde{Y}_{q}$, $Y_{l}$, $\tilde{Y}_{l}$, $Y_{\Delta_{\mathrm{L}}}$, and $Y_{\Delta_{\mathrm{R}}}$ represent the Yukawa coupling matrices. The dilutor $\Delta$ is the Higgs boson arising from the electrically neutral component of $\Delta_{\mathrm{R}}$. Its decay channels involve $\Delta\rightarrow N_{1}N_{1}$, $\Delta\rightarrow W_{\mathrm{R}}^{\ast}W_{\mathrm{R}}^{\ast}\to\text{4 fermions}$, $\Delta\rightarrow\gamma\gamma$, etc.
@@
We solve the Boltzmann equations for the benchmark points (BPs) specified in Tables 1 and 2 for Scenarios 1 and 2, respectively. The values of $y$, $m_{N_{1}}$, $m_{N_{2}}$, $m_{\Delta}$, $m_{W_{\mathrm{R}}}$, the ratio $\tan\beta$ of two vacuum expectation values from the bidoublet, and the mixing angle $\theta_{\Delta h}$ between $\Delta$ and the $125~$\mathrm{G}\mathrm{e}\mathrm{V}$$ Higgs boson $h$ are adopted based on the results presented in Ref. [68]. The dilutor decay width $\Gamma_{Y}$ and branching ratio $B_{X}$ are calculated using the formulas provided therein. All BPs are selected to ensure the correct DM relic abundance through the DM dilution mechanism. BP1a and BP2a will be the primary focus of the following analysis, while the remaining BPs will provide additional results for comparison.
We adopt the initial temperature as $T_{\mathrm{ini}}=m_{Y}/10$, at which $Y$ particles have decoupled with $Y_{Y}\simeq Y_{X}$ and are nonrelativistic, while $X$ particles are relativistic. The initial conditions for the energy densities of $Y$, $X$, and SM particles are
@@
By substituting the obtained $\rho_{Y}$, $\rho_{X}$, and $\rho_{\mathrm{SM}}$ into Eq. (45) and solving Eq. (9) with $a(t_{0})=1$, we determine the evolution of the scale factor $a(t)$ in both BPs, as illustrated in Fig. 4. Compared with the standard $\Lambda\mathrm{CDM}$ cosmological model, the presence of the EMD era reduces the scale factor prior to $t_{2}$. This occurs because the scale factor evolves as $a\propto t^{2/3}$ during an MD era, increasing more rapidly with time than $a\propto t^{1/2}$ during an RD era. Therefore, a smaller $a$ is required at the onset of the EMD era to ensure $a(t_{0})=1$ at the present time $t_{0}$.
@@ III.3 Impact on the loop number density of cosmic strings
Since the insertion of the EMD era affects the evolution of the CS network in the VOS model, it is necessary to examine how the normalized correlation length $\xi$ and RMS velocity $v$ change with time. We should consider these variations when calculating the number density of CS loops during the EMD era. To achieve this, we solve Eqs. (16) and (17) with the Hubble rate $H(t)$ modified by the presence of the EMD era and obtain $\xi$ and $v$ as functions of cosmic time $t$ for BP1a and BP2a, as demonstrated by blue solid lines in Fig. 5. For comparison, the results of the standard $\Lambda\mathrm{CDM}$ model without the EMD era are presented as red dotted lines.
@@
During the RD era in the $\Lambda\mathrm{CDM}$ model, $\xi$ and $v$ maintain the scaling values $\xi_{\mathrm{r}}=0.271$ and $v_{\mathrm{r}}=0.662$, as stated in Eq. (18). Some deviations around $t\sim 10^{20}~$\mathrm{G}\mathrm{e}\mathrm{V}^{-1}$$ and $t\sim 10^{26}~$\mathrm{G}\mathrm{e}\mathrm{V}^{-1}$$ are attributed to changes in relativistic degrees of freedom shown in Eq. (12). For $t\gtrsim t_{\mathrm{eq}}$, the universe transits to the MD era, resulting in an increase in $\xi$ and a decrease in $v$. It is important to note that the time axis in Fig. 5 is on a logarithmic scale, indicating that the changes in $\xi$ and $v$ occur relatively slowly. Nonetheless, this represents a mild nonscaling effect.
When the EMD era occurs, $\xi$ increases, while $v$ decreases, and both quantities gradually return to their scaling values after the EMD era ends. Therefore, the EMD era also introduces a nonscaling effect, which will be incorporated into the calculation of the CS loop number density. Furthermore, we estimate the magnitudes of the time derivatives $\dot{\xi}$ and $\dot{v}$ by calculating the evolution of their ratios to $\xi/t$ and $v/t$, respectively, as shown in Fig. 6. We find that during the EMD era for both BPs, $|t\dot{\xi}/\xi|$ and $|t\dot{v}/v|$ are smaller than 0.150.150.15 and 0.030.030.03, respectively. Thus, $\dot{\xi}$ and $\dot{v}$ are negligible compared to $\xi/t$ and $v/t$, simplifying the calculations below.
@@
where $\Gamma G\mu\ll\alpha_{\mathrm{m}}\xi_{\star}=0.18$ is used at the third step.
In addition, similar to Eq. (35), we must account for the number densities of CS loops produced in a preceding era and surviving into subsequent eras following the timeline in Fig. 2. For clarity, we define the notations for these loop number densities as follows.
@@
which is equivalent to Eq. (35). Similarly, by introducing $l_{1,2}=l+\Gamma G\mu(t-t_{1,2})$, we obtain
@@ III.4 Modification of the GW spectrum
To assess the sensitivity of GW detection experiments, we also show the constraints from the North American Nanohertz Observatory for Gravitational Waves (NANOGrav) [96], European Pulsar Timing Array (EPTA) [97], and Parkes Pulsar Timing Array (PPTA) [98], as well as the sensitivity curves of the International Pulsar Timing Array (IPTA) [99], Square Kilometer Array (SKA) [100], Laser Interferometer Space Antenna (LISA) [52], TianQin [101], Taiji [102], LIGO [103], Cosmic Explorer (CE) [104], and Einstein Telescope (ET) [105].
@@
Figure 7: GW spectra from cosmic strings with $G\mu=10^{-11}$ modified by the EMD era for BP1a in Scenario 1 (a) and BP2a in Scenario 2 (b). The red solid lines represent the total GW spectra, which include contributions from CS loops across all eras. The blue, green, olive, and cyan dotted lines correspond to the contributions from CS loops formed during the original RD, EMD, second RD, and final MD eras, respectively. For comparison, various constraints and sensitivity curves from several GW detection experiments are also shown.
Compared to the GW spectrum for $G\mu=10^{-11}$ in the $\Lambda\mathrm{CDM}$ model illustrated in Fig. 1, the spectra in both BPs with the EMD era display a suppression at high frequencies $f\gtrsim 10~$\mathrm{H}\mathrm{z}$$, which corresponds to the contributions from CS loops formed during the original RD and EMD eras. The primary reason for this suppression is as follows. Since the scaling behavior of the CS network is only slightly violated, the correlation length $L$ remains approximately proportional to the cosmic time $t$, suggesting that the lengths of the generated loops are positively correlated with the scale factor $a(t)$. As shown in Fig. 4, the EMD era reduces the scale factor before $t_{2}$. This means that CS loops with a given initial length $l$, which is inversely proportional to the frequencies of the emitted GWs through $f_{\mathrm{e}}=2n/l$, are formed at a later time when the energy densities of both CS loops and emitted GWs are reduced. Consequently, the GW spectrum at sufficiently high frequencies is suppressed, and the end time of the EMD era, $t_{2}$, is related to the GW frequency at which this suppression effect becomes significant. For the parameters listed in Table 1, the dilutor $N_{2}$ in BP1a has a longer lifetime than the dilutor $\Delta$ in BP2a, resulting in a longer EMD era and, consequently, a stronger suppression.
@@
In addition, we consider the results for BP1b and BP1c in Scenario 1 and for BP2b and BP2c in Scenario 2, whose parameters are tabulated in Tables 1 and 2, respectively. The obtained GW spectra are illustrated in Fig. 9. For BP1a, BP1b, and BP1c in Scenario 1, the dilutor $N_{2}$ has the same mass but the masses of the DM candidate $N_{1}$ differ, causing variations in the decay width of $N_{2}$. A smaller decay width $\Gamma_{N_{2}}$ corresponds to a longer duration of the EMD era, leading to stronger suppression effects at high frequencies, as illustrated in Fig. 9(a). In contrast, the masses of the dilutor $\Delta$ are different for BP2a, BP2b, and BP2c in Scenario 2. A heavier $\Delta$ implies that the EMD era occurs earlier, and hence, a higher frequency at which the GW spectrum begins to be suppressed. This is clearly demonstrated in Fig. 9(b).
@@
where $[f_{\mathrm{min}},f_{\mathrm{max}}]$ is the accessible frequency band. If the SNR reaches a threshold $\varrho_{\mathrm{thr}}=10$, it is probable that a GW signal will be detected.
@@
For $G\mu\lesssim 10^{-12}$, the inclusion of the EMD era has an insignificant impact on the SNRs because the suppression effect occurs at frequencies outside the sensitive bands of LIGO, CE, and ET (cf. Fig. 8). In contrast, for $G\mu\gtrsim 10^{-12}$, the SNRs are obviously reduced. Compared to BP2a, the dilutor in BP1a has a longer lifetime, leading to a stronger suppression effect and, consequently, smaller SNRs.